# Forward and Backward State Abstractions for Off-policy Evaluation

### v1

@@ D.2 Proof of Theorem 2
    $V_{1}^{\pi}(s^{(1)})=$ $\mathbb{E}^{\pi}\left[R_{1}|S_{1}=s^{(1)},A_{1}=a\right]\pi(a|s^{(1)})$
    $=$ $\mathcal{R}(a,s^{(1)})\pi(a|s^{(1)})$
    $=$ $\underbrace{\mathcal{R}(a,s^{(2)})}_{\mbox{reward-irrelevant}}\underbrace{\pi(a|s^{(2)})}_{\pi-\mbox{irrelevant}}$
@@
    $Q_{j}^{\pi}(a,s^{(1)})=Q_{j}^{\pi}(a,s^{(2)})\,\,\mbox{and}\,\,V_{j}^{\pi}(s^{(1)})=V_{j}^{\pi}(s^{(2)}).$ (D.3)
    Our goal is to show (D.3) holds with $j=T$.
    We similarly define $Q_{j,\phi}^{\pi}$ and $V_{j,\phi}^{\pi}$ as the Q- and value functions defined on the abstract state space. Similar to the proof of Theorem 1, we can show that $Q_{j}^{\pi}=Q_{j,\phi}^{\pi}$ and $V_{j}^{\pi}=V_{j,\phi}^{\pi}$ for any $j<T$. It follows that
@@
    $=$ $\gamma\mathbb{E}^{\pi}\sum_{s^{\prime}\in\mathcal{S}}\left[\sum_{t=2}^{T}\gamma^{t-1}R_{t}|S_{2}=s^{\prime}\right]\mathcal{T}(s^{\prime}|s^{(1)},a)+\mathcal{R}(a,s^{(1)})$
@@
[…]
    $=$ $Q^{\pi}_{T}(a,s^{(2)}).$
    This together with $\pi$-irrelevance proves $V_{T}^{\pi}$-irrelevance. Consequently, (D.3) holds for any $j\geq 1$. Since $Q_{j}^{\pi}\to Q^{\pi}$ as $j\to\infty$, we obtain $Q^{\pi}$-irrelevance.
  • We will prove that the MIS estimator constructed on the abstract state space remains valid. With a slight abuse of notation, we use $p_{t}^{\pi}(a,x)$ to denote the probability $\mathbb{P}^{\pi}(A_{t}=a,\phi(S_{t})=x)$. Under the stationarity assumption, direct calculations yield
@@
    $=$ $\mathbb{E}\left[(1-\gamma)^{-1}w_{\phi}^{\pi}(A,\phi(S))\underbrace{\mathcal{R}\big{(}A,\phi(S)\big{)}}_{\mbox{reward-irrelevant}}\right]$
@@
    $=$ $\mathbb{E}[f_{3}(w^{\pi})]$
    Notice that we only require reward-irrelevance in the above proof.
  • It suffices to show that
    $\mathbb{E}[\rho_{1:t}^{\pi}R_{t}]=\mathbb{E}[\prod_{k=1}^{t}\rho_{\phi,t}^{\pi}(A_{k},\phi(S_{k}))R_{t}],$ (D.4)
@@
    Using similar arguments in proving the validity of MIS estimator, under reward-irrelevance, (D.5) can be shown to equal to
    $\sum_{a\in\mathcal{A},x\in\mathcal{X}}p_{t}^{\pi}(a,x)\mathcal{R}_{\phi}(a,x).$ (D.6)
    Under transition-irrelevance, the data triplets $(\phi(S),A,R)$ forms an MDP, satisfying the Markov assumption. Let $\mathcal{T}_{\phi}$ denote the resulting transition function. Together with $\pi$-irrelevance, we can rewrite (D.6) as
    $\sum_{\begin{subarray}{c}a_{1},\cdots,a_{t}\in\mathcal{A}\\ x_{1},\cdots,x_{t}\in\mathcal{X}\end{subarray}}\rho_{0}(x_{1})\prod_{k=1}^{t-1}\Big{[}\pi_{\phi}(a_{k}|x_{k})\mathcal{T}_{\phi}(x_{k+1}|a_{k},x_{k})\Big{]}\pi_{\phi}(a|x_{t})\mathcal{R}_{\phi}(a,x).$
    Notice that $\mathcal{T}_{\phi}$ is independent of the target policy $\pi$. Using the change of measure theorem, we can represent above expression by $\mathbb{E}(\rho_{1:t,\phi}^{\pi}R_{t})$ where $\rho_{1:t,\phi}^{\pi}$ denotes the cumulative IS ratio defined on the abstract state space. This completes the proof.
  • Since model-irrelvance implies $Q^{\pi}$-irrelevance, the conclusion directly follows from the last conclusion of Theorem 1.
@@ D.3 Proof of Theorem 3
At the begging of the proof, we name the phenomena as the Inverse Markovianity, namely the reversed state-action pairs maintain the Markov property.
  • $\rho^{\pi}$-irrelevance directly follows from the definition of backward-model-irrelevance. To show $w^{\pi}$-irrelevance, we divide the proof into two steps.
    (1) In the first step, we will prove that if $\phi$ satisfies the backward-model-irrelevance, then
    $\rho^{\pi}(A_{t-k},S_{t-k})\perp\!\!\!\perp S_{t}|\phi(S_{t}),1\leq k\leq t-1.$ (D.7)
    It follows from equation (8) that
    $\mathbb{P}\big{(}\phi(S_{t-k})=x|S_{t-k+1}\big{)}=\mathbb{P}\big{(}\phi(S_{t-k})=x|\phi(S_{t-k+1})\big{)},1\leq k\leq t-1.$
    We can use the induction method to prove that for $11\leq k\leq t-$ 1,
    $\rho^{\pi}(A_{t-k},S_{t-k})\perp\!\!\!\perp S_{t}|\phi(S_{t}).$ (D.8)
    For $k=1$, we have for any positive constant $c$,
    $\mathbb{P}\big{(}\rho^{\pi}(A_{t-1},S_{t-1})=c|S_{t}\big{)}=$ $\mathbb{P}[\rho_{\phi,t-1}^{\pi}\big{(}A_{t-1},\phi(S_{t-1})\big{)}=c|S_{t}]$
    $=$ $\mathbb{P}[\rho_{\phi,t-1}^{\pi}\big{(}A_{t-1},\phi(S_{t-1})\big{)}=c|\phi(S_{t})],$ (D.9)

### v2

@@ D.2 Proof of Lemma 1
    $V_{1}^{\pi}(s^{(1)})=$ $\sum_{a\in\mathcal{A}}Q_{1}^{\pi}(a,s^{(1)})\pi(a|s^{(1)})$
    $=$ $\sum_{a\in\mathcal{A}}Q_{1}^{\pi}(a,s^{(2)})\pi(a|s^{(2)})$
@@
    $Q_{j}^{\pi}(a,s^{(1)})=Q_{j}^{\pi}(a,s^{(2)})\,\,\mbox{and}\,\,V_{j}^{\pi}(s^{(1)})=V_{j}^{\pi}(s^{(2)}),$ (D.4)
    whenever $\phi(s^{(1)})=\phi(s^{(2)})$. Our goal is to show (D.4) holds with $j=T$.
    We similarly define $Q_{j,\phi}^{\pi}$ and $V_{j,\phi}^{\pi}$ as the Q- and value functions on the abstract state space. Similar to the proof of Theorem D.1, we can show that $Q_{j}^{\pi}=Q_{j,\phi}^{\pi}$ and $V_{j}^{\pi}=V_{j,\phi}^{\pi}$ for any $j<T$.
    Direct calculations yield
@@
    $=$ $\mathbb{E}^{\pi}\sum_{s^{\prime}\in\mathcal{S}}\left[\sum_{t=2}^{T}\gamma^{t-1}R_{t}|S_{2}=s^{\prime}\right]\mathcal{T}(s^{\prime}|s^{(1)},a)+\mathcal{R}(a,s^{(1)})$
@@
    $=$ $Q^{\pi}_{T}(a,s^{(2)}),$
    where the second last equation holds due to transition-irrelevance in (3) and (D.4), which states that $V_{T-1}^{\pi}(s^{\prime})$ is constant as a function of $s^{\prime}\in\phi^{-1}(x^{\prime})$.
    This together with $\pi$-irrelevance proves $V_{T}^{\pi}$-irrelevance. By induction, we have shown that (D.4) holds for any $j\geq 1$. Under the boundness assumption in Assumption 1, $Q_{j}^{\pi}\to Q^{\pi}$ as $j\to\infty$. We thus obtain (D.3), which yields $Q^{\pi}$-irrelevance.
    Next, we prove the identifiability of $Q^{\pi}_{\phi}$. Similarly, we define
    $Q_{j,\phi}^{\pi}(a,x)=\sum_{t=1}^{j}\gamma^{t-1}\mathbb{E}^{\pi}\left[R_{t}|\phi(S_{1})=x,A_{1}=a\right].$ (D.5)
    By setting $j=1$, it reduces to $\mathbb{E}^{\pi}[R_{1}|\phi(S_{1})=x,A_{1}=a]$. Under the MDP model assumption, the conditional mean of the immediate reward depends solely on the current state-action pair, independent of past history. This together with the reward-irrelevance condition further implies that the conditional mean of the reward depends solely on the abstract-state-action pair. Consequently,
    $\mathbb{E}^{\pi}[R_{1}|\phi(S_{1})=x,A_{1}=a]=\underbrace{\mathbb{E}[R_{1}|\phi(S_{1})=x,A_{1}=a]}_{\mathcal{R}_{\phi}(a,x)}.$
    Notice that the expectation on the right-hand-side (RHS) is defined with respect to the behavior policy. It can thus be consistently estimated using the offline data under the coverage assumption in Assumption 2. This yields the identifiability of $Q_{1,\phi}^{\pi}$.
    Similarly, we can show that
    $\mathbb{P}^{\pi}[\phi(S_{2})=x^{\prime}|A_{1}=a,\phi(S_{1})=x]=\underbrace{\mathbb{P}[\phi(S_{2})=x^{\prime}|A_{1}=a,\phi(S_{1})=x]}_{\mathcal{T}_{\phi}(x^{\prime}|a,x)},$
    under transition-irrelevance, which establishes the identifiability of the left-hand-side (LHS).
    Notice that for any $j\geq 1$, under the MDP model assumption, $Q_{j,\phi}^{\pi}$ can be represented using $\mathbb{E}^{\pi}[R_{1}|\phi(S_{1})=x,A_{1}=a]$ and $\mathbb{P}^{\pi}[\phi(S_{2})=x^{\prime}|A_{1}=a,\phi(S_{1})=x]$. Both have been proven identifiable. This the establishes identifiability of $Q_{j,\phi}^{\pi}$. Again, by letting $j\to\infty$, we obtain the identifiability of $Q_{\phi}^{\pi}$ under the boundedness assumption in Assumption 1. The proof is hence completed.
  • Fisher consistency of MIS. We use $p_{t,\phi}^{\pi}(a,x)$ to denote the probability $\mathbb{P}^{\pi}(A_{t}=a,\phi(S_{t})=x)$ and $p_{t}^{\pi}(s)$ to denote $\mathbb{P}^{\pi}(S_{t}=s)$. Under the stationary assumption, direct calculations yield
@@
    $=$ $\mathbb{E}\left[(1-\gamma)^{-1}w_{\phi}^{\pi}(A,\phi(S))\underbrace{\mathcal{R}_{\phi}\big{(}A,\phi(S)\big{)}}_{\mbox{reward-irrelevant}}\right]$
@@
    $=$ $\mathbb{E}[f_{3}(w^{\pi})].$
    To complete the proof, it remains to establish the identifiability of $w_{\phi}^{\pi}$.
    Under the stationarity assumption in Assumption 3, $\omega_{\phi}^{\pi}$ can be represented by
    $\frac{\sum_{t\geq 1}\gamma^{t-1}p_{t,\phi}^{\pi}(a,x)}{\mathbb{P}(A_{1}=a,\phi(S_{1})=x)}.$
    It is immediate to see that the denominator is identifiable, as the probability is calculated with respect to the behavior policy. It suffices to show that for any $t\geq 1$, $p_{t}^{\pi}$ is identifiable as well. Under transition-irrelevance, the data triplets $(\phi(S),A,R)$ forms an MDP, satisfying the Markov assumption. As such, we can rewrite $p_{t}^{\pi}(a_{t},x_{t})$ as
    $\sum_{\begin{subarray}{c}a_{1},\cdots,a_{t-1}\in\mathcal{A}\\ x_{1},\cdots,x_{t-1}\in\mathcal{X}\end{subarray}}\rho_{0,\phi}(x_{1})\prod_{k=1}^{t-1}\Big{[}\pi_{\phi}(a_{k}|x_{k})\mathcal{T}_{\phi}(x_{k+1}|a_{k},x_{k})\Big{]}\pi_{\phi}(a_{t}|x_{t}),$
    where $\rho_{0,\phi}$ denotes the pmf of $\phi(S_{1})$ which is independent of $\pi$, and both $\pi_{\phi}$ and $\mathcal{T}_{\phi}$ are identifiable under $\pi$- and transition-irrelevance, respectively. This proves the identifiability of $p_{t}^{\pi}$, and hence, the identifiability of $w_{\phi}^{\pi}$.
  • Fisher consistency of SIS. Recall that we require the behavior policy to be Markovian, meaning that at any time $t$, $A_{t}$ is independent of historical observations given $S_{t}$. A key challenge in state abstraction for the SIS estimator is that, after abstraction, the behavior policy can be history-dependent, leading to the inconsistency of SIS. Toward that end, we employ a history-dependent IS ratio to address this challenge. Specifically, let $\rho_{j,\phi}^{\pi}$ denote
    $\rho_{j,\phi}^{\pi}=\frac{\pi_{\phi}(A_{j}|\phi(S_{j}))}{b_{j,\phi}(A_{j}|\phi(S_{j}),A_{j-1},\phi(S_{j-1}),\ldots,A_{1},\phi(S_{1}))}.$ (D.6)
    Under $\pi$-irrelevance, the numerator is well-defined and identifiable. However, unlike the standard IS ratio where the denominator depends solely on the current state, the denominator in (D.6) depends on the entire history. Notice that the denominator is essentially the data distribution of $A_{j}$ given $\phi(S_{j}),A_{j-1},\phi(S_{j-1}),\ldots,A_{1},\phi(S_{1})$, it is thus identifiable from the offline data as well.
    Under the coverage assumption in Assumption 2, the behavior policy is bounded away from zero. Since the behavior policy is stationary, this conditional pmf can be represented by
    $\mathbb{E}[b(\bullet|S_{j})|\phi(S_{j}),A_{j-1},\phi(S_{j-1}),\ldots,A_{1},\phi(S_{1})],$
    which is bounded away from zero as well. Consequently, $\rho_{t,\phi}^{\pi}$ is bounded and identifiable.
    Let $\rho_{1:t,\phi}^{\pi}$ denote the SIS ratio $\prod_{j=1}^{t}\rho_{j,\phi}^{\pi}$. It suffices to show
    $\mathbb{E}(\rho_{1:t}^{\pi}R_{t})=\mathbb{E}(\rho_{1:t,\phi}^{\pi}R_{t}),$ (D.7)
@@
    Using similar arguments to the proof of Fisher consistency of MIS estimator, under reward-irrelevance, (D.8) can be shown to equal to
    $\sum_{\begin{subarray}{c}a_{1},\cdots,a_{t}\in\mathcal{A}\\ x_{1},\cdots,x_{t}\in\mathcal{X}\end{subarray}}\rho_{0}(x_{1})\prod_{k=1}^{t-1}\Big{[}\pi_{\phi}(a_{k}|x_{k})\mathcal{T}_{\phi}(x_{k+1}|a_{k},x_{k})\Big{]}\pi_{\phi}(a_{t}|x_{t})\mathcal{R}_{\phi}(a_{t},x_{t}).$
    Notice that both $\mathcal{T}_{\phi}$ and $\mathcal{R}_{\phi}$ independent of the target policy $\pi$. Using the change of measure theorem, we can represent above expression by $\mathbb{E}(\rho_{1:t,\phi}^{\pi}R_{t})$. This completes the proof.
  • Fisher consistency of DRL under model-irrelevance. Since model-irrelevance and $\pi$-irrelevance imply $Q^{\pi}$-irrelevance and the identifiability of $Q^{\pi}$, the conclusion directly follows from the double robustness of DRL and that in the first bullet point.
@@ D.3 Proof of Theorem 1
We establish the Fisher consistencies of SIS, MIS, Q-function-based method and DRL one by one.
  • Fisher consistency of SIS. Notice that $\rho^{\pi}$-irrelevance directly follows from the definition of backward-model-irrelevance and $\pi$-irrelevance. It follows from Lemma D.1 that $\mathbb{E}[f_{2}(\rho^{\pi})]=\mathbb{E}[f_{2}(\rho^{\pi}_{\phi})]$.
    Additionally, under $\pi$-irrelevance and behavior-policy-irrelevance, both the numerator and the denominator of the IS ratio $\rho_{\phi}^{\pi}$ are identifiable. Consequently, $\rho_{\phi}^{\pi}$ is identifiable as well. This establishes the Fisher consistency of SIS.
  • Fisher consistency of MIS. We first establish the $w^{\pi}$-irrelevance. We next establish the identifiability of $w_{\phi}^{\pi}$.
    To prove the $w^{\pi}$-irrelevance, we begin by defining the marginal density ratio at a given time $t$ as
    $w^{\pi}_{t}(a,s)=\frac{\mathbb{P}^{\pi}(A_{t}=a,S_{t}=s)}{\mathbb{P}(A_{t}=a,S_{t}=s)}.$
    Under the stationarity assumption, the denominator is independent of $t$. Notice that $w^{\pi}(a,s)=(1-\gamma)\sum_{t=1}^{\infty}\gamma^{t-1}w_{t}^{\pi}(a,s)$. Hence, it is sufficient to prove that $\phi$ is $w_{t}^{\pi}$-irrelevance, for any $t$.
    We prove this by induction. First, when $t=1$, $w_{t}^{\pi}$ is reduced to $\rho^{\pi}$. Consequently, $w_{1}^{\pi}$-irrelevance is readily obtained by backward-model-irrelevance and $\pi$-irrelevance.
    Second, suppose we have established $w_{t}^{\pi}$-irrelevance. We aim to show $w_{t+1}^{\pi}$-irrelevance. With some calculations, we obtain that
    $\begin{split}&\mathbb{E}[w^{\pi}_{t}(A_{t},S_{t})\rho^{\pi}(A_{t+1},S_{t+1})|A_{t+1}=a,S_{t+1}=s]\\ =&\sum_{a^{\prime},s^{\prime}}w^{\pi}_{t}(a^{\prime},s^{\prime})\frac{\pi(a|s)}{b(a|s)}\mathbb{P}(A_{t}=a^{\prime},S_{t}=s^{\prime}|A_{t+1}=a,S_{t+1}=s)\\ =&\sum_{a^{\prime},s^{\prime}}\frac{\mathbb{P}^{\pi}(A_{t}=a^{\prime},S_{t}=s^{\prime})}{\mathbb{P}(A_{t}=a^{\prime},S_{t}=s^{\prime})}\frac{\pi(a|s)}{b(a|s)}\frac{\mathbb{P}(A_{t}=a^{\prime},S_{t}=s^{\prime},A_{t+1}=a,S_{t+1}=s)}{\mathbb{P}(A_{t+1}=a,S_{t+1}=s)}\\ =&\sum_{a^{\prime},s^{\prime}}\mathbb{P}^{\pi}(A_{t}=a^{\prime},S_{t}=s^{\prime})\frac{\pi(a|s)}{b(a|s)}\times\frac{\mathbb{P}(A_{t+1}=a|S_{t+1}=s,A_{t}=a^{\prime},S_{t}=s^{\prime})}{\mathbb{P}(A_{t+1}=a,S_{t+1}=s)}\mathbb{P}(S_{t+1}=s|A_{t}=a^{\prime},S_{t}=s)\\ =&\sum_{a^{\prime},s^{\prime}}\mathbb{P}^{\pi}(A_{t}=a^{\prime},S_{t}=s^{\prime})\frac{\pi(a|s)}{\mathbb{P}(A_{t+1}=a,S_{t+1}=s)}\mathbb{P}(S_{t+1}=s|A_{t}=a^{\prime},S_{t}=s)\\ =&\frac{\mathbb{P}^{\pi}(A_{t+1}=a,S_{t+1}=s)}{\mathbb{P}(A_{t+1}=a,S_{t+1}=s)}\\ =&w^{\pi}_{t+1}(a,s),\end{split}$ (D.9)
    where the third last equality is due to that the behavior policy is stationary. This establishes the link between $w_{t}^{\pi}$, $\rho^{\pi}$ and $w_{t+1}^{\pi}$.
    To prove $w_{t+1}^{\pi}$-irrelevance, we first prove the following equation holds:
    $\begin{split}&\sum_{s^{\prime}\in\phi^{-1}(x^{\prime})}\mathbb{P}(A_{t}=a^{\prime},S_{t}=s^{\prime}|A_{t+1}=a,S_{t+1}=s^{(1)})\\ =&\sum_{s^{\prime}\in\phi^{-1}(x^{\prime})}\mathbb{P}(A_{t}=a^{\prime},S_{t}=s^{\prime}|A_{t+1}=a,S_{t+1}=s^{(2)}),\end{split}$ (D.10)
    whenever $\phi(s^{(1)})=\phi(s^{(2)})$.
    Indeed, by (5), we obtain that
    $\sum_{s^{\prime}\in\phi^{-1}(x^{\prime})}\mathbb{P}(A_{t}=a^{\prime},S_{t}=s^{\prime}|A_{t+1}=a,S_{t+1}=s^{(1)})$
    $=$ $\sum_{s^{\prime}\in\phi^{-1}(x^{\prime})}\frac{\mathbb{P}(A_{t}=a^{\prime},S_{t}=s^{\prime},A_{t+1}=a,S_{t+1}=s^{(1)})}{\mathbb{P}(A_{t+1}=a,S_{t+1}=s^{(1)})}$
    $=$ $\sum_{s^{\prime}\in\phi^{-1}(x^{\prime})}\frac{\mathbb{P}(A_{t+1}=a|S_{t+1}=s^{(1)},A_{t}=a^{\prime},S_{t}=s^{\prime})}{\mathbb{P}(A_{t+1}=a,S_{t+1}=s^{(1)})}\mathbb{P}(S_{t+1}=s^{(1)},A_{t}=a^{\prime},S_{t}=s^{\prime})$
    $=$ $\sum_{s^{\prime}\in\phi^{-1}(x^{\prime})}\mathbb{P}(A_{t}=a^{\prime},S_{t}=s^{\prime}|S_{t+1}=s^{(1)})$
    $=$ $\sum_{s^{\prime}\in\phi^{-1}(x^{\prime})}\mathbb{P}(A_{t}=a^{\prime},S_{t}=s^{\prime}|S_{t+1}=s^{(2)})$
    $=$ $\sum_{s^{\prime}\in\phi^{-1}(x^{\prime})}\mathbb{P}(A_{t}=a^{\prime},S_{t}=s^{\prime}|A_{t+1}=a,S_{t+1}=s^{(2)}).$
    Consequently,
    $w^{\pi}_{t+1}(a,s^{(1)})$
    $=$ $\sum_{a^{\prime},s^{\prime}}w^{\pi}_{t}(a^{\prime},s^{\prime})\rho^{\pi}(a,s^{(1)})\mathbb{P}(A_{t}=a^{\prime},S_{t}=s^{\prime}|A_{t+1}=a,S_{t+1}=s^{(1)})$
    $=$ $\sum_{a^{\prime},s^{\prime}}w^{\pi}_{t}(a^{\prime},s^{\prime})\rho^{\pi}(a,s^{(2)})\mathbb{P}(A_{t}=a^{\prime},S_{t}=s^{\prime}|A_{t+1}=a,S_{t+1}=s^{(2)})$
    $=$ $w^{\pi}_{t+1}(a,s^{(2)}),$